A rectangular plate 10ft long and 6ft wide is submerged vertically with the shorter side 2ft below the water surface. Find the force on one side of the plate


1
Expert's answer
2021-05-25T18:45:00-0400

There will be a hydrostatic pressure acting on a system when the body is submerged in liquid ,


Hydrostatic pressure at depth h = "{\rho}gh"


where "{\rho}=" density of water =1000 kg/m"^{3}"


g = acceleration due to gravity = 9.8 m/s"^{2}"


Total force = Pressure * area



"=\int_{0}^{10}\int_{0}^{6}{\rho}gh\ dh\ dx"



"={\rho}g\dfrac{[h^{2}]}{2}_{0}^{6}" "\int_{0}^{10}dx"


"=18{\rho}g(10-0)" =180 "\times 1000\times9.8=" 1764000"\ N"